If the difference between the real depth and apparent depth of a swimming pool filled with water is 1.6m, determine the depth of the pool Assume refractive index of water is 1.33


1
Expert's answer
2020-08-24T13:11:58-0400

"Real \space depth-Apparent\space depth=1.6"


"Apparent \space depth=\frac{real\space depth}{refractive\space index\space of\space water}"


Assume that apparent depth= x therefore, real depth= apparent depth(x)+1.6


substituting this in the above equation


"x=\frac{x+1.6}{1.33}"


"1.33x-x=1.6"


"0.33x=1.6"


"x=4.848"


Therefore Apparent depth="4.848"


The real depth of the pool is therefore "4.848+1.6=6.448m"


Answer= 6.448m